OpenStudy (anonymous):

please help me... I know the answer is -30. I just need the steps: 42=-3(m+6)=m

OpenStudy (anonymous):

something wrong here 42 = etc. = -m means m=-42

OpenStudy (anonymous):

42=-3(m+6)+m.....Sorry

OpenStudy (anonymous):

42 = -3m -18 + m = 2m - 18 60 = 2 m you are nearly done

OpenStudy (anonymous):

Then divid it by 2

OpenStudy (nikato):

*42=-3m-18+m +18. +18 ------------------- 60= -2m --- ---- -2. -2 m=?

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

Thxs